# Optimal ambition in business, politics and life

### v3

@@ II.3 Social comparison hurts performance and penalizes ambition
Fig. 4 shows simulations for 100-agent cohorts. Agents vary in their satisfaction thresholds, each chosen randomly from a uniform distribution. The optimal satisfaction threshold is slightly lower with cohort comparison than when agents know the true reward distribution. Because agents can become unsatisfied if the cohort statistics change, setting a lower threshold prevents agents from overreacting to others’ success. Additional cohort sizes and threshold intervals are shown in Fig. A1 and Fig. A2.
@@
Figure 4: Social comparison is costly. Mean rewards for cohort comparisons and upward social comparisons are shown on a rough, non-skewed landscape. Cohort comparison (yellow curve) lowers the optimal satisfaction threshold and cumulative reward. Upward social comparison (purple curve) further lowers the optimal threshold and substantially hinders performance. Both the expected reward and the optimal satisfaction threshold are lower compared to the reference landscape (black, dashed curve), which assumes no social comparison. Parameter values are: $\epsilon_{t}\sim\mathcal{N}(\mu,\sigma^{2})$, $\varphi=0$, and $t_{\text{max}}=1000$. Cohorts include $100$ agents, with thresholds uniformly sampled from $[-3,6]$. Results shown are averaged over $10^{4}$ simulations.
II.4 Applications
@@
Fig. 5 illustrates examples of where our model’s insights regarding smoothness and skewness could be applied to economic policy, entrepreneurship, and political campaigns. These examples reveal a subtle but important distinction between optimal ambition and optimal risk-taking.
Economic growth rates tend to have left-skewed distributions in developed countries, because recessions tend to have larger magnitudes than booms Burgess et al. (2020b). Fig. 5a shows this pattern in U.S. data. From a policy standpoint, the left-skewed distribution of economic growth rates implies that risk taking is costly. If downside risks are larger than upside risks, risk-taking lowers the expected outcome. However, our model implies seemingly opposite advice for ambition, compared to the mean reward. A country’s mean economic growth rate is dragged down by large and rare recessions (Fig. 5a). Policymakers would therefore be unwise to be satisfied with growth rates close to this mean in a typical year. In other words, the left-skewness of economic growth rates calls for less risky but more ambitious (compared to the mean) economic policymaking.
Private wealth distributions have the opposite property: they are heavily right-skewed Scheffer et al. (2017). Fig. 5b illustrates the wealth distributions of the world’s richest billionaires and the world’s largest companies. A right-skewed distribution of private wealth outcomes implies that upside risks are often larger than downside risks—risk-taking in entrepreneurship often means either winning big or losing small. Consequently, risk taking is often beneficial in business and private investment Benuzzi and Ploner (2024). However, our model suggests that the right skewness of private wealth calls for less ambition, compared to the mean available reward. This is counterintuitive, but is related to the fact that mean success in wealth building and entrepreneurship is highly influenced by rare and extremely successful ‘unicorns’. Thus, an entrepreneur or investor who sets a high satisfaction threshold compared to the mean may miss opportunities for important and achievable successes.
Fig. 5c shows the skewness and smoothness of monthly polling margins in swing states ahead of the 2020 U.S. presidential election Five Thirty Eight (2020), from the perspective of the Republican party. (For the Democrats, skewness would have the opposite sign and smoothness would be the same.) Higher margins are more likely in states with more rugged and left-skewed polling distributions. Our model implies that parties should pursue more ambitious (albeit less risky, given the left-skew) political strategies in such states.
For example, pre-election polling distributions in 2020 would have suggested that Republicans had a higher chance of outperforming their mean polling margins in Minnesota and South Carolina than they did in neighboring Wisconsin and Georgia (Fig. 5c). This suggests that the Republicans had a greater incentive to pursue ambitious campaign strategies in these states, all else equal. Of course, there are many other factors, besides expected achievable vote margins, that are relevant in allocating campaign resources to states in practice. Such factors include (but are not limited to) the Electoral College rewards and how close the expected vote margins are to zero, given the first-past-the-post system in most states.
@@
Figure 5: Panel a: U.S. growth rates in gross domestic product (GDP) per capita (gray) Our World in Data (2024), compared moving means (black) and medians (dashed). The medians are greater than the means, indicating left skew. Panel b: distributions of private wealth among the 2500 top companies Dhiman (2024) and billionaires Elgiriyewithana (2023). The means are much greater than the medians, indicating right skew. Panel c: monthly polling margins, from the perspective of Republicans (i.e. Republican share -- Democrat share) in U.S. swing states during 2020, prior to the presidential election Five Thirty Eight (2020). Darker colors indicate larger average differences between the two parties. Lighter-colored states have smaller average margins.
III Discussion
@@
Our model then adds more precise insights. First, we show that being too unsatisfiable is costlier than being too easily satisfied, on average (Fig. 2a). Second, we show that left-skewed and rugged reward distributions should motivate more ambition (Fig. 3). This exposes a subtle difference between optimal ambition and optimal risk taking, as left-skewed rewards motivate less risk-taking Benuzzi and Ploner (2024). Third, we show that upward social comparison is costly by creating unrealistic expectations (Fig. 4). We describe how these insights apply to real-world contexts relating to economic policy, wealth building, and political campaigns (Fig. 5). We discuss additional real-world examples in the introduction.
@@
Our model is admittedly simple; indeed, its accessibility to a broad audience is one of its contributions. However, ‘explore-exploit’ and sequential searches are common abstractions in areas such as microeconomics Kohn and Shavell (1974), management science Lavie et al. (2010); Greve (2007), harvesting Richard et al. (2018), and animal foraging MacArthur and Pianka (1966); Cowie (1977). Our ‘satisfaction’ framework is somewhat related to the satisficing concept in behavioral economics Simon (1955), but it is also distinct in that our threshold is optimized, rather than being a behavioral heuristic.
@@
The framework of optimal search has been previously studied through different lenses in different disciplines, some of which provide insights into how our results might extend to more complex settings. For example, Kohn and Shavell Kohn and Shavell (1974) consider how time preference, risk aversion, uncertainty, or search costs affect optimal search. Their model is too complex for optimal satisfaction thresholds to be derived in general, but they show that risk aversion, search costs, and discounting future rewards would each reduce optimal satisfaction thresholds. These insights are intuitive. They also show that uncertainty can raise optimal satisfaction thresholds by increasing the information value of searching Kohn and Shavell (1974). This finding is somewhat counterintuitive in light of our finding that higher-than-optimal thresholds are costlier than lower-than-optimal thresholds (Fig. 2a), combined with that fact that left-skewed uncertainty typically motivates risk avoidance Benuzzi and Ploner (2024).
@@
Applied mathematicians have widely studied what they call ‘the secretary problem’, where one seeks to maximize the chance of selecting the best option from a sequentially observed set Freeman (1983). For example, MacQueen and Miller MacQueen and Miller Jr (1960) characterize the optimal search duration under these conditions. The exact bounds and optimal strategies depend on the underlying distribution and corresponding assumptions. In the case of AR1 processes (like our model), several bounds have been derived for exponentially distributed steps  Novikov (2009); Christensen (2012); Christensen et al. (2011). However, because of their complexities and sensitivities to assumptions, empirical applications often rely on powerful but opaque machine learning approaches Becker et al. (2019); Ciocan and Mišić (2022).
@@
Social comparisons have been theoretically predicted and empirically documented to affect decision-making in numerous human and animal contexts. Humans evaluate their successes and status relative to their peers Layard (2006, 2011), and, as a result, they make consumption decisions aimed at optimizing both absolute and relative outcomes (e.g., ref. Arrow and Dasgupta (2009)). Human fact-finding and decision-making processes can be either enhanced (e.g., the wisdom of crowds Surowiecki (2005)) or corrupted (e.g., groupthink Janis (1971)) by social forces. In animals, instincts for social conformism have also been linked to improved (e.g., ref. Berdahl et al. (2013)) and diminished (e.g., ref. Dussutour et al. (2007)) decision-making in different contexts.
@@
Future research should explore other complexities that our model does not consider. For example, there can be inherent benefits in reality to sticking with one strategy (e.g., accruing expertise, deepening a relationship). Similarly to search costs (e.g., disruptions from moving cities or changing jobs), this would likely reduce optimal satisfaction thresholds. Real people are also prone to weighing potential losses more than potential gains (known as ‘loss aversion’ Kahneman et al. (1991)). This, too, would reduce the optimal satisfaction threshold in absolute terms in our model, by reducing the utility that agents assign to low-end nominal rewards. Our agents are simplified in terms of both their decisions and their inferences about available rewards. Our social comparison model does not consider structured social networks. We model random rather than intelligent search processes, though it is straightforward to conjecture that intelligent search would increase optimal ambition by making high but rare rewards easier to find. There are many other worthwhile extensions to our framework.
@@
Despite these limitations, our conceptual results are intuitive and likely to generalize. For example, increasing ambition makes satisfactory rewards harder to find under any search process. Making high rewards more abundant—via skewness or ruggedness—makes ambitious targets easier to achieve. Upward social comparison distorts perceptions of reality, which hinders rational decision making.
@@
How ambitious to be is an important question in a wide range of professional, political, and personal contexts. Our model, results, and examples offer entry points for mathematically precise, but also accessible and intuitive, explorations of this question.
IV Methods
We summarize our main analytical proofs below. Additional details and the simulation algorithm can be found in the appendix.
IV.1 The expected reward distribution in a rugged, Gaussian landscape
With a rugged landscape (i.e. $\varphi=0$) with a Gaussian reward distribution $\epsilon_{t}\sim\mathcal{N}(\mu=0,\sigma^{2})$, with PDF $\phi(x)$ and CDF $\Phi(x)$, we compute the expected reward by thinking of the satisfaction threshold as a divider between the ‘explore’ phase and the ‘exploit’ phase. The rewards collected during the ‘explore’ phase are, by definition, less than the threshold value, so that $X_{\text{explore}}<T$. The rewards collected during the ‘exploit’ phase are greater than or equal to the threshold value: $X_{\text{exploit}}\geq T$. Rewards earned during the exploration phase have the lower-tail truncated normal distribution with the mean $\mu_{\text{explore}}$, and the single reward which is repeatedly collected during the exploitation phase has the upper-tail truncated normal distribution with the mean $\mu_{\text{exploit}}$.
For threshold $T$, the mean $\mu_{\text{explore}}=-\dfrac{\phi(T)}{\Phi(T)}$ is given by the inverse Mills ratio Greene (2003). The agents expect to do worse than the landscape mean in the exploration phase, on average ($\mu_{\text{explore}}<0$). The expected reward earned during the exploitation phase is given by $\mu_{\text{exploit}}=\dfrac{\phi(T)}{1-\Phi(T)}>0$.
$\mathbb{E}[\text{reward}]$ $=\sum_{t_{x}=0}^{t_{x}=t_{\text{max}}}\Bigg{[}\mu_{\text{explore}}(T)t_{x}+\mu_{\text{exploit}}(T)\left(t_{\text{max}}-t_{x}\right)\Bigg{]}$
×$\quad\times(1-\Phi(T))\Phi(T)^{t_{x}}.$ (2)
To standardize our results across different time spans ($t_{\text{max}}$), we normalize the cumulative reward by dividing by $t_{\text{max}}$ and compare expected reward per time step (e.g., Fig. 2).
IV.2 Effect of smoothness on the expected reward distribution
For a landscape with smoothness $\varphi$, the variance of the associated AR1 process is $\text{var}[X_{t}]=(1+\varphi)/(1-\varphi)$. The analytical expression for the expected cumulative reward on smooth landscapes can therefore be derived by simply scaling the threshold $T$ by a factor of $1/\sqrt{\text{var}[X_{t}]}$ in Equation 2.
IV.3 The optimal satisfaction threshold is finite and increases as the total time increases
For any threshold, $T$, the probability of finding a strategy that satisfies the threshold increases with the length of the search, $t_{\text{max}}$. The amount of time one gets to exploit a satisfactory strategy also increases in $t_{\text{max}}$, all else equal. Both of these patterns shift incentive towards a larger optimal threshold as $t_{\text{max}}$ increases. In the limiting case of $t_{\text{max}}\rightarrow\infty$, the expected reward per time step is equal to $\mu_{\text{exploit}}$ (see Appendix A for details), meaning that when time is infinite, agents can be arbitrarily ambitious, but optimal ambition is finite in finite time.
IV.4 The optimal satisfaction threshold is always greater than the mean reward
The reward in the ‘exploit’ phase ($\mu_{\text{exploit}}(T)$) is always positive and increases in $T$. The expected number of time steps it takes to satisfy threshold $T$ is $\frac{1}{1-\Phi(T)}$: at $T$=0, the expected length of the exploration phase is two time steps, and grows exponentially as threshold $T$ increases. Intuitively then, agents who can afford the exploration phase to be longer than two time steps should target positive thresholds $T$.
We analytically show that the optimal threshold is above the mean for the maximally rugged landscape. We can split the summand in  (2) into the reward component
$\mu_{\text{explore}}(T)t_{x}+\mu_{\text{exploit}}(T)\left(t_{\text{max}}-t_{x}\right),$
and the probability component,
$(1-\Phi(T))\Phi(T)^{t}.$
For a given $t_{x}$, the reward component increases near-linearly in $T$.
The probability component is unimodal and
achieves a single maximum when $\Phi(T)=\frac{t_{x}}{t_{x}+1}$.
Note that the value of $T$ that maximizes the probability component increases in $t_{x}$ and is above zero for $t_{x}>0$. The probability component decays exponentially in $T$ around the maximum, as well as in $t_{x}$. We can think of the probability component as a filter on the reward component. When we pointwise multiply an increasing function with a unimodal hump-shaped function, the resulting maximum of the summand occurs for a threshold that satisfies $T\geq\Phi^{-1}(\frac{t_{x}}{t_{x}+1})$.
Consider now (2), the sum over all possible time steps $t_{x}$ where the satisfaction thresholds could potentially be met. Summing over $t_{x}$, we are summing over a sequence of functions in $T$ which are hump-shaped and whose maxima occur at higher thresholds $T$ for higher time step value $t_{x}$. Therefore the optimal threshold $T$ for the sum lies for $T>\Phi^{-1}(\frac{1}{2})=0$.
Since the reward component increases linearly in $T$, and the probability component decreases exponentially in $T$, the contribution from each subsequent $t_{x}$ term is decreasing, and hence the optimal satisfaction threshold is finite. We elaborate on this argument in the appendix.
IV.5 Data Availability
For part of this work, we used previously published data Bruch and Newman (2018); Hoxby and Avery (2012); Five Thirty Eight (2020); Our World in Data (2024); Dhiman (2024); Elgiriyewithana (2023). Simulation code is available on GitHub at https://github.com/kathlandgren/ambition.
Appendix A: Derivation and limiting behavior of expected reward

### v4

@@ III.4 Social comparison hurts performance and penalizes ambition
Fig. 4 shows simulations for 100-agent cohorts. Agents vary in their satisfaction thresholds, each chosen randomly from a uniform distribution. The optimal satisfaction threshold is slightly lower with cohort comparison than when agents know the true reward distribution. Because agents can become unsatisfied if the cohort statistics change, setting a lower threshold prevents agents from overreacting to others’ success. Additional cohort sizes and threshold intervals are shown in Fig. 10 and Fig. 11 in Appendix C.
@@
Figure 4: Social comparison is costly. Mean rewards for cohort comparisons and upward social comparisons are shown on a rough, non-skewed landscape. Cohort comparison (yellow curve) lowers the optimal satisfaction threshold and cumulative reward. Upward social comparison (purple curve) further lowers the optimal threshold and substantially hinders performance. Both the expected reward and the optimal satisfaction threshold are lower compared to the reference landscape (black, dashed curve), which assumes no social comparison. Parameter values are: $\epsilon_{t}\sim\mathcal{N}(0,1)$, $\varphi=0$, and $t_{\text{max}}=1000$. Cohorts include 100100 agents, with thresholds uniformly sampled from [3,6][-3,6]. Results shown are averaged over $10^{4}$ simulations.
III.5 Search costs
To consider the effects of search costs, we assume that the agent pays a constant cost, $c$, in each period they spend searching for a new strategy. Inserting this assumption into equation (1), the agent’s reward in time period $t$, $X_{t}$, becomes:
$X_{t}=\begin{cases}\varphi X_{t-1}+(1-\varphi)\epsilon_{t}-c&X_{t-1}<T,\\ X_{t-1}&X_{t-1}\geq T.\end{cases}$ (3)
From equation (2), their expected reward becomes:
$\mathbb{E}[\text{reward}]$ $=$
$\sum_{t_{x}=0}^{t_{\text{max}}}\Bigg[(\mu_{\text{explore}}(T)-c)t_{x}+\mu_{\text{exploit}}(T)\left(t_{\text{max}}-t_{x}\right)\Bigg]$
$\times(1-\Phi(T))\Phi(T)^{t_{x}}.$ (4)
Figure 5: Effects of search costs. Expected rewards for search with a per-timestep search cost on a rough, non-skewed landscape. The cost values are: (a) $c=2$, (b) $c=10$, (c) $c=395$, (d) $c=800$. Introducing a search cost (colored curves) lowers the optimal satisfaction threshold and can produce negative expected rewards when either the cost or the threshold are too high. In all cases, both the expected reward and the optimal satisfaction threshold are lower than those of the reference landscape (black dashed curve). The sums are computed numerically based on equation (4). Parameter values are: $\epsilon_{t}\sim\mathcal{N}(0,1)$, $\varphi=0$, and $t_{\text{max}}=1000$.
Fig. 5 shows the effects of increasing the search costs ($c$) on the expected reward as a function of the satisfaction threshold ($T$). Search costs gradually reduce both the optimal satisfaction threshold and the maximum achievable reward (Fig. 5a). Search costs increase the degree to which being too ambitious is costlier than being too cautious, such that being too ambitious can now make expected rewards worse than the expected rewards of never searching (i.e., $\mu$ per time step) (Fig. 5b). As long as a satisfaction threshold exists with higher expected rewards than never searching, however small, the optimal satisfaction threshold will be still be higher than the mean of available rewards ($\mu$) (Fig. 5c). Otherwise, the optimal threshold will be -$\infty$ (i.e., the agent should never search) (Fig. 5d).
In other words, search costs only affect our main qualitative result—that optimal ambition targets finite rewards strictly larger than the mean of available rewards—in the extreme case where costs are so high that they prohibit all search.
III.6 Applications
@@
Although our model is stylized and does not literally describe real-world decision-making in its full complexity, Fig. 6 illustrates examples of where our model’s qualitative insights regarding smoothness and skewness could be applied to economic policy, entrepreneurship and political campaigns. These examples reveal a subtle but important distinction between optimal ambition and optimal risk-taking, which should generalize beyond our stylized setting. It follows directly from the statistical properties of the reward distribution, and not from the specific search procedure assumed.
@@
Figure 6: Possible applications. Panel a: U.S. growth rates in gross domestic product (GDP) per capita (gray) [40], compared moving means (black) and medians (dashed). The medians are greater than the means, indicating left skew. Panel b: distributions of private wealth among the 2500 top companies [41] and billionaires [42]. The means are much greater than the medians, indicating right skew. Panel c: monthly polling margins, from the perspective of Republicans (i.e. Republican share - Democrat share) in U.S. swing states during 2020, prior to the presidential election [43]. Darker colors indicate larger average differences between the two parties. Lighter-colored states have smaller average margins.
Economic growth rates tend to have left-skewed distributions in developed countries, because recessions tend to have larger magnitudes than booms [44]. Fig. 6a shows this pattern in U.S. data. From a policy standpoint, the left-skewed distribution of economic growth rates implies that risk taking is costly. If downside risks are larger than upside risks, risk-taking lowers the expected outcome. However, our model implies seemingly opposite advice for ambition, compared to the mean reward. A country’s mean economic growth rate is dragged down by large and rare recessions (Fig. 6a). Policymakers would therefore be unwise to be satisfied with growth rates close to this mean in a typical year. In other words, the left-skewness of economic growth rates calls for less risky but more ambitious (compared to the mean) economic policymaking.
Private wealth distributions have the opposite property: they are heavily right-skewed [45]. Fig. 6b illustrates the wealth distributions of the world’s richest billionaires and the world’s largest companies. A right-skewed distribution of private wealth outcomes implies that upside risks are often larger than downside risks—risk-taking in entrepreneurship often means either winning big or losing small. Consequently, risk taking is often beneficial in business and private investment [34]. However, our model suggests that the right skewness of private wealth calls for less ambition, compared to the mean available reward. This is counterintuitive, but it is related to the fact that mean success in wealth building and entrepreneurship is highly influenced by rare and extremely successful ‘unicorns’. Thus, an entrepreneur or investor who sets a high satisfaction threshold compared to the mean may miss opportunities for important and achievable successes.
Fig. 6c shows the skewness and smoothness of monthly polling margins in swing states ahead of the 2020 U.S. presidential election [43], from the perspective of the Republican party. (For the Democrats, skewness would have the opposite sign and smoothness would be the same.) Parties and candidates undertake platforms, messaging, and on-the-ground campaign strategies that either explicitly or implicitly aim for certain vote margin targets in different states and regions. Higher margins are more likely in states with more rugged and left-skewed polling distributions. Therefore, our model implies that parties should pursue more ambitious (albeit less risky, given the left-skew) political strategies in such states, all else being equal.
For example, pre-election polling distributions in 2020 would have suggested that Republicans had a higher chance of outperforming their mean polling margins in Minnesota and South Carolina than they did in neighboring Wisconsin and Georgia (Fig. 6c). This suggests that the Republicans had a greater incentive to pursue ambitious campaign strategies in these states, all else equal. Of course, there are many other factors, besides expected achievable vote margins, that are relevant in allocating campaign resources to states in practice. Such factors include (but are not limited to) the Electoral College rewards, tradeoffs and synergies between strategies targeting different regions and constituencies, and how close the expected vote margins are to zero, given the first-past-the-post system in most states.
IV Discussion
@@
Our model then adds more precise insights. First, we show that being too unsatisfiable is costlier than being too easily satisfied, on average (Fig. 2a). Second, we show that left-skewed and rugged reward distributions should motivate more ambition (Fig. 3). This exposes a subtle difference between optimal ambition and optimal risk taking, as left-skewed rewards motivate less risk-taking [34]. Third, we show that upward social comparison is costly by creating unrealistic expectations (Fig. 4). Fourth, we show that our model’s main insight is insensitive to search costs, up the point where search costs disincentivize all search (Fig.  5). We describe how these insights apply qualitatively to real-world contexts relating to economic policy, wealth building and political campaigns (Fig. 6). We discuss additional real-world examples in the introduction.
@@
Our model is admittedly simple; indeed, its accessibility to a broad audience is one of its contributions. However, ‘explore-exploit’ and sequential searches are common abstractions in areas such as microeconomics [15], management science [46, 47], harvesting [48] and animal foraging [49, 21]. Our ‘satisfaction’ framework is somewhat related to the satisficing concept in behavioral economics [50], but it is also distinct in that our threshold is optimized, rather than being a behavioral heuristic.
@@
The framework of optimal search has been previously studied through different lenses in different disciplines, some of which provide insights into how our results might extend to more complex settings. For example, Kohn and Shavell [15] consider how time preference, risk aversion and uncertainty affect optimal search. Their model is too complex for optimal satisfaction thresholds to be derived in general, but they show that risk aversion and discounting future rewards would each reduce optimal satisfaction thresholds. These insights are intuitive. They also show that uncertainty can raise optimal satisfaction thresholds by increasing the information value of searching [15]. This finding is somewhat counterintuitive in light of our finding that higher-than-optimal thresholds are costlier than lower-than-optimal thresholds (Fig. 2a), combined with that fact that left-skewed uncertainty typically motivates risk avoidance [34].
@@
Applied mathematicians have widely studied ‘the secretary problem’, where one seeks to maximize the chance of selecting the best option from a sequentially observed set [13]. For example, MacQueen and Miller [11] characterize the optimal search duration under these conditions. The exact bounds and optimal strategies depend on the underlying distribution and corresponding assumptions. In the case of AR(1) processes (like our model), several bounds have been derived for exponentially distributed steps  [25, 26, 27]. However, because of their complexities and sensitivities to assumptions, empirical applications often rely on powerful but opaque machine learning approaches [24, 16].
@@
Social comparisons have been theoretically predicted and empirically documented to affect decision-making in numerous human and animal contexts. Humans evaluate their successes and status relative to their peers [51, 52], and, as a result, they make consumption decisions aimed at optimizing both absolute and relative outcomes (e.g., ref. [53]). Human fact-finding and decision-making processes can be either enhanced (e.g., the wisdom of crowds [54]) or corrupted (e.g., groupthink [55]) by social forces. In animals, instincts for social conformism have also been linked to improved (e.g., ref. [56]) and diminished (e.g., ref. [57]) decision-making in different contexts.
@@
Future research should explore other complexities that our model does not consider. For example, there can be inherent benefits in reality to sticking with one strategy (e.g., accruing expertise, deepening a relationship). Similarly to search costs (e.g., disruptions from moving cities or changing jobs), this would likely reduce optimal satisfaction thresholds. Real people are also prone to weighing potential losses more than potential gains (known as ‘loss aversion’ [58]). This, too, would reduce the optimal satisfaction threshold in absolute terms in our model, by reducing the utility that agents assign to low-end nominal rewards. Our agents are simplified in terms of both their decisions and their inferences about available rewards. Our social comparison model does not consider structured social networks. We model random rather than intelligent search processes, though it is straightforward to conjecture that intelligent search would increase optimal ambition by making high but rare rewards easier to find. There are many other worthwhile extensions to our framework.
@@
Despite these limitations, our conceptual results are intuitive and likely to generalize. For example, increasing ambition makes satisfactory rewards harder to find under any search process. Making high rewards more abundant—via skewness or ruggedness—makes ambitious targets easier to achieve. Upward social comparison distorts perceptions of reality, which hinders rational decision making. Search costs reduce optimal ambition, but optimal ambition still generally exceeds the mean of available rewards.
@@
How ambitious to be is an important question in a wide range of professional, political and personal contexts. Our model, results and examples offer entry points for mathematically precise, but also accessible and intuitive, explorations of this question.
Data availability
For parts of this work, we used previously published data, available from refs. [31, 32, 43, 40, 41, 42].
Code availability
Our simulation code is available on GitHub at https://github.com/kathlandgren/ambition. The simulation algorithm is described in Appendix B.
Acknowledgements
We thank members of the Burgess lab, Ellen DeGennaro, Todd Cherry, Jonas L. Juul, Chase Thiel, and Scott Beaulier for comments. E.L. acknowledges funding from the Cooperative Institute for Research in Environmental Sciences (CIRES) Visiting Fellows Program at the University of Colorado Boulder, funded by NOAA Cooperative Agreement NA22OAR4320151.
Appendix A Additional derivations
A.1 Limiting behavior as $t_{\text{max}}\rightarrow\infty$